nP6=6(nP4)
so far I have: n!/(n-6)!=6[n!/(n-4)!]

Answer by scott8148(6628)

dividing by n! and multiplying by (n-4)! ___ (n-4)! / (n-6)! = 6

(n-4)(n-5) = 6 ___ n^2 - 9n + 14 = 0

factoring ___ (n-7)(n-2) = 0